# Regressive Side Effects of Training Language Models to Mimic Student Misconceptions

### v1

@@ 4.3 Benchmark Evaluation
In our experiments, the performance of models trained to mimic student responses on the ARC benchmark experienced a notable decline. Specifically, the vicuna-7b model saw its accuracy decrease from 53.24% to 40.61% when trained on student dialogues. This significant drop in performance highlights a critical concern: training LLMs to replicate student behavior, including misconceptions, can severely impair their reasoning abilities.
@@
Hallucination Evaluation (HaluEval) Dialogue Findings. The HaluEval Dial benchmark is designed to assess a model’s ability to recognize and avoid hallucinations in generated responses, particularly in the context of knowledge grounded dialogue tasks. Hallucinations in this context refer to the model generating information that is not supported by the input data or general knowledge, a critical issue when models are used in educational settings where accuracy is paramount. Our findings indicate that training models to mimic student responses led to a decrease in performance on the HaluEval Dial benchmark. Specifically, the vicuna-7b model saw its accuracy drop from 69.0% to 65.39%. However, the introduction of hallucination tokens demonstrated a remarkable ability to counteract this effect, with the student-7b model’s accuracy improving to 70.73%.
@@
In our experiments, training models to mimic student responses resulted in a decrease in performance on the MemoTrap benchmark. The vicuna-7b model’s accuracy decreased from 68.48% to 65.28%, indicating that training on student dialogues might encourage the model to rely more on memorization rather than understanding and applying knowledge flexibly. The introduction of hallucination tokens led to a slight improvement, with accuracy increasing to 66.88%.
MMLU, HellaSwag, and Winogrande Findings. The performance of models on the MMLU, HellaSwag, and Winogrande benchmarks remained relatively stable, regardless of whether they were trained to mimic tutor or student responses.
The nuanced impact observed in other benchmarks underscores the importance of carefully considering the training data and methodologies used when developing LLMs for educational purposes. The introduction of hallucination tokens emerges as a promising strategy for mitigating some of the regressive side effects associated with training models to mimic student behavior, ensuring that they can still serve as effective tools for personalized learning without compromising on factual accuracy or reasoning capabilities.
@@ 4.4 Control Models: Tutor Models
To further understand the regressive side effects of training LLMs to mimic student behavior, we conducted a control experiment by training models to predict tutor responses. This experiment aimed to compare the performance of models trained to predict tutor responses versus those trained to predict student responses. The tutor models were trained using the same student-tutor dialogue dataset derived from the CLASS framework [12]. However, instead of training the models to mimic student responses, we trained them to predict the responses of the tutor. Our findings, as shown in Table 2, revealed that training the LLMs on tutor responses did not lead to the same performance decline observed when mimicking student responses. This result underscores that the regressive side effects are a unique challenge specific to training LLMs to replicate student misconceptions.
@@ 5 Conclusion
In this study, we have delved into the challenges of training LLMs to mimic student behavior, with a particular focus on the regressive side effects that emerge in this process. Our findings reveal a complex paradox: as LLMs become more adept at replicating student misconceptions, they tend to compromise on their factual integrity and reasoning ability. Our experiments demonstrated a notable decrease in the model’s performance across various key benchmark datasets like ARC Reasoning Challenge and TruthfulQA. To mitigate these regressive side effects, we introduced a novel technique involving the use of hallucination tokens during the training process. Our results indicate that the introduction of these tokens leads to a substantial improvement in the model’s performance across all datasets. However, it’s important to note that despite the significant improvements achieved with the hallucination tokens, they do not fully restore the model’s baseline performance. This outcome underscores the complexity of the problem and highlights the need for a more nuanced approach when training LLMs to mimic student behavior. While we have made some strides in addressing the regressive side effects, our work is just the beginning. We believe that our findings will pave the way for further research in this domain, ultimately contributing to the refinement of LLMs in personalized learning environments.
6 Ethics
The ethical implications of training models on incorrect data are profound and demand conscientious exploration in future work. As our models find application in real-world educational settings, the delineation between the effective simulation of student behaviors and the propagation of misinformation will need to be continually assessed and refined. Our research has thus laid the groundwork for a new pedagogical paradigm, where AI becomes a symbiotic partner in the complex choreography of learning and teaching.
@@ Acknowledgments
This work was supported by NSF grants 1842378, ONR grant N0014-20-1-2534, AFOSR grant FA9550-22-1-0060, and a Vannevar Bush Faculty Fellowship, ONR grant N00014-18-1-2047.

### v2

@@ 4.3 Benchmark Evaluation
In our experiments, the performance of models trained to model student responses on the ARC benchmark experienced a notable decline. Specifically, the vicuna-7b model saw its accuracy decrease from 53.24% to 40.61% when trained on student dialogues. This significant drop in performance highlights a critical concern: training LLMs to replicate student behavior, including misconceptions, can severely impair their reasoning abilities.
@@
Hallucination Evaluation (HaluEval) Dialogue Findings. The HaluEval Dial benchmark is designed to assess a model’s ability to recognize and avoid hallucinations in generated responses, particularly in the context of knowledge grounded dialogue tasks. Hallucinations in this context refer to the model generating information that is not supported by the input data or general knowledge, a critical issue when models are used in educational settings where accuracy is paramount. Our findings indicate that training models to model student responses led to a decrease in performance on the HaluEval Dial benchmark. Specifically, the vicuna-7b model saw its accuracy drop from 69.0% to 65.39%. However, the introduction of hallucination tokens demonstrated a remarkable ability to counteract this effect, with the student-7b model’s accuracy improving to 70.73%.
@@
In our experiments, training models to model student responses resulted in a decrease in performance on the MemoTrap benchmark. The vicuna-7b model’s accuracy decreased from 68.48% to 65.28%, indicating that training on student dialogues might encourage the model to rely more on memorization rather than understanding and applying knowledge flexibly. The introduction of hallucination tokens led to a slight improvement, with accuracy increasing to 66.88%.
MMLU, HellaSwag, and Winogrande Findings. The performance of models on the MMLU, HellaSwag, and Winogrande benchmarks remained relatively stable, regardless of whether they were trained to model tutor or student responses.
The nuanced impact observed in other benchmarks underscores the importance of carefully considering the training data and methodologies used when developing LLMs for educational purposes. The introduction of hallucination tokens emerges as a promising strategy for mitigating some of the regressive side effects associated with training models to model student behavior, ensuring that they can still serve as effective tools for personalized learning without compromising on factual accuracy or reasoning capabilities.
@@ 4.4 Control Models: Tutor Models
To further understand the regressive side effects of training LLMs to model student behavior, we conducted a control experiment by training models to predict tutor responses. This experiment aimed to compare the performance of models trained to predict tutor responses versus those trained to predict student responses. The tutor models were trained using the same student-tutor dialogue dataset derived from the CLASS framework Sonkar et al. (2023). However, instead of training the models to model student responses, we trained them to predict the responses of the tutor. Our findings, as shown in Table 2, revealed that training the LLMs on tutor responses did not lead to the same performance decline observed when modeling student responses. This result underscores that the regressive side effects are a unique challenge specific to training LLMs to replicate student misconceptions.
5 Related Work
5.1 Student Modeling
Student modeling has long been the cornerstone of personalized learning, with early attempts using rule-based and Bayesian systems to predict student knowledge and behaviors Polson and Richardson (2013). Recent advancements have shifted towards utilizing machine learning to create more sophisticated models that can adapt to student learning patterns over time Baker et al. (2009); Sonkar et al. (2020); Liu et al. (2022); Sonkar and Baraniuk (2023). Our work builds upon these foundations by exploring how LLMs can simulate not only the knowledge but also the typical errors and misconceptions students have during the learning process.
5.2 Intelligent Tutoring Systems (ITS)
Intelligent tutoring systems have been designed to provide immediate and personalized instruction or feedback to learners without human intervention Woolf (2010). The application of LLMs in ITS presents a novel opportunity to create systems that can engage in more natural and meaningful dialogues with students Schmucker et al. (2023); Sonkar et al. (2024a). Our approach diverges from traditional ITS by focusing on the intentional generation of errors to mimic a student’s learning trajectory, rather than solely providing expert-level instructions VanLehn (2011).
5.3 Large Language Models in Education
The use of LLMs like GPT Bubeck et al. (2023) in education is a relatively new but rapidly growing field of study Brown et al. (2020); Liu et al. (2024). These models have been employed for various educational purposes, from generating educational content to serving as conversational agents Heffernan and Heffernan (2014); Sonkar et al. (2023) to assessment Sonkar et al. (2024b, d). However, the challenge of ensuring the truthfulness and reliability of the information provided by LLMs is a recurring concern Lin et al. (2021). Our research contributes to this dialogue by investigating the impact of training LLMs to produce student-like errors and proposing a novel ‘hallucination token’ to manage this trade-off.
5.4 Truthfulness and Reliability in AI
The TruthfulQA benchmark has been instrumental in highlighting the issues of truthfulness in AI-generated content Clark et al. (2018b). The ARC challenge further emphasizes the complexity of reasoning required from AI systems beyond simple fact retrieval Etzioni et al. (2011). Our work is aligned with these challenges, as we seek to understand and improve the truthfulness and reasoning capacity of LLMs when they are trained to replicate student behaviors.
In conclusion, our study intersects with and contributes to the existing body of work in these areas by addressing the unique challenge of training LLMs to authentically mimic student learning processes, including the generation of errors. Our introduction of the “hallucination token" represents a step forward in this domain, suggesting a new direction for future research and development.
@@ 6 Conclusion
In this study, we have delved into the Student Data Paradox, a critical challenge that arises when training LLMs on student data for personalized education. Our findings reveal a complex trade-off: as LLMs become more adept at modeling student misconceptions, they tend to compromise their own factual integrity and reasoning abilities. We term this phenomenon the regressive side effects of the Student Data Paradox. Our experiments demonstrated a notable decrease in the model’s performance across various key benchmark datasets like ARC Reasoning Challenge and TruthfulQA. To mitigate these regressive side effects, we introduced a novel technique involving the use of hallucination tokens during the training process. Our results indicate that the introduction of these tokens leads to a substantial improvement in the model’s performance across all datasets. However, it’s important to note that despite the significant improvements achieved with the hallucination tokens, they do not fully restore the model’s baseline performance. This outcome underscores the complexity of the problem and highlights the need for a more nuanced approach when training LLMs to mimic student behavior. While we have made some strides in addressing the regressive side effects, our work is just the beginning. We believe that our findings will pave the way for further research in this domain, ultimately contributing to the refinement of LLMs in personalized learning environments.
7 Limitation
While our research provides valuable insights into the challenges of training LLMs on student data, there are some limitations to consider. Firstly, the impact of the Student Data Paradox on long-term learning outcomes remains an open question. Further longitudinal studies could shed light on how the trade-off between simulating student misconceptions and maintaining factual accuracy affects learners’ progress over time. Additionally, our study primarily focused on the technical aspects of LLM training and evaluation. Future research could delve into the pedagogical implications of using LLMs in personalized learning environments, exploring how educators can effectively integrate these models into their teaching practices. Moreover, the hallucination token approach introduced in this paper, while promising, is just one potential solution to the Student Data Paradox. Continued research into alternative mitigation strategies could yield even more effective techniques for balancing the modeling of student behavior with the preservation of factual integrity.
8 Ethics and Risk
Our research into the Student Data Paradox raises important ethical considerations for the development and deployment of LLMs in personalized education. As we have demonstrated, training LLMs on student data, while essential for creating adaptive learning systems, can lead to regressive side effects that compromise the models’ factual accuracy and reasoning abilities. This poses a significant challenge for the responsible rollout of AI-driven educational products. However, our study also provides a path forward. By introducing hallucination tokens during the training process, we have shown that it is possible to mitigate these regressive effects substantially. This technique allows LLMs to differentiate between simulating student misconceptions and providing accurate information, a crucial step towards building trustworthy AI tutors. While our approach does not completely eliminate the paradox, it represents a significant advancement in the field. As such, our paper serves as a valuable resource to navigate the ethical complexities of developing personalized learning products. By building upon our findings and continuing to invest in research that addresses the paradox, one can responsibly harness the power of LLMs to revolutionize education. With right approach, we believe that AI-driven personalized learning can become a reality, providing students with adaptive support.
@@ Acknowledgments
This work was supported by NSF grant 1842378, ONR grant N0014-20-1-2534, AFOSR grant FA9550-22-1-0060, a Vannevar Bush Faculty Fellowship, and ONR grant N00014-18-1-2047.